# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v1

@@ 5.1 The model
The recent conceptual and technical advances in fMRI data analysis provide an opportunity for the field to not only the effectiveness of brain disease diagnosis deepen our understanding of brain disease mechanisms but also deepen our understanding of brain disease mechanisms (Zheng et al., 2024; Luo et al., 2025; Kan et al., 2022b; Yu et al., 2024; Wang et al., 2022; Hao et al., 2023; Bian et al., 2024; Bhattacharya et al., 2025). However, current methods overlook more higher-order patterns with signs, limiting an integrated understanding of brain-wide communication for brain disease diagnosis.
@@
In this study, we focused on the importance of signed higher-order(group) interactions in fMRI signals, which were inferred using a recently developed topological approach (Hyekyoung et al., 2011). Our proposed HOI-Brain incorporates: (i) a novel higher-order interaction representation strategy via Multiplication of Temporal Derivatives (MTD) to quantify dynamic functional co-fluctuations of group ROIs; (ii) a novel high-order feature extraction method that links higher-order interactions (HOIs) in the brain to higher-order organizational structures, focusing on quadruplet-level interaction signatures and two-dimensional void descriptors extracted through Persistent Homology; (iii) a pioneering effort to distinguish between positively and negatively synergistic higher-order interactions in the brain; (iv) a novel multi-channel brain network Transformer that integrates lower-order edge features with higher-order topological invariants. This new framework emphasizes the potential of quadruplet or higher-order region interactions—representing critical advancements for precision medicine, a deeper understanding of neurological disorders, and broader contributions to neuroimaging research. It demonstrates superior prediction accuracy for classifying Alzheimer’s Disease (AD), Parkinson’s Disease (PD), and Autism Spectrum Disorder (ASD) compared to traditional machine learning models, graph neural network (GNN)-based models, transformer-based models, hypergraph neural network (HGNN)-based models, and the Persistent Homology (PH)-based models. And ablation study also indicates that the effectiveness of quadruplet interactions compared to triplet interactions, the effectiveness of distinguishing between positively and negatively synergistic higher-order interactions, and the effectiveness of different components in model.
@@
Our investigations revealed intriguing patterns. When designing an orthonormal clustering readout to identify clusters of functionally similar nodes through soft clustering with orthonormal constraints, we found that assigning nodes from different modes to distinct functional modules significantly improves classification accuracy. This suggests that both low-level and high-level feature patterns—often residing within distinct functional modular structures—play a critical role, This phenomenon is consistent with findings reported in previous literature (Wang et al., 2019). This new discovery also provides a new opportunity for brain disease diagnostic models—specifically, how to effectively identify and integrate low-order and high-order patterns across different functional modules in the brain.
@@
In addition, our investigation into the most discriminative neural patterns using attention mechanism revealed that quadruplet-level interaction signatures exhibit substantially greater importance than two-dimensional void descriptors and edge features in both the ADNI and ABIDE datasets. Notably, negative synergistic quadruplets generally demonstrate higher importance than their positive synergistic quadruplets, whereas positive synergistic voids surpass negative synergistic voids in significance. This novel observation may offer critical insights into the pathological mechanisms underlying brain disorders. To further validate these findings, we performed inter-group comparisons of signed high-order topological signatures in the ADNI dataset. The results revealed highly significant group differences in the number of quadruplet-level interaction signatures (all $p<0.001$ ) compared to void descriptors). Moreover, positive voids exhibited more pronounced inter-group differences than negative voids. These findings suggest that disease progression may primarily disrupt higher-order neural interactions, which subsequently influence high-dimensional hole structures. An intriguing pattern emerged in the comparative analysis of cognitively normal (CN) individuals and those with Alzheimer’s disease (AD). CN individuals appear to rely more on positive information for whole-brain topological organization, whereas AD patients show a greater dependence on negative information. Furthermore, longitudinal trends indicate that the number of positive quadruplets gradually declines with disease progression, while negative quadruplets follow a biphasic trajectory—decreasing initially and then increasing. This implies that during the transition from CN to mild cognitive impairment (MCI), both positive and negative synergistic interactions diminish. However, as the condition advances from MCI to AD, neural networks increasingly adopt negative synergy while suppressing excitability, possibly as a compensatory mechanism to preserve functional integrity across neural modules.
@@
When exploring the important brain regions and their interactions associated with brain diseases, we use HOI-Brain to identify them from a more comprehensive perspective, rather than the low-level perspective adopted in previous studies (Yan et al., 2019; Li et al., 2021). These findings, serving as potential biomarkers, are consistent with a large body of existing literature reports (Zidan et al., 2019; Zhi et al., 2024; Ye et al., 2019; Kim et al., 2021), and also provide new insights for brain disease diagnosis. Specifically, the right caudate may serve as a hub node within the brain network and exhibits disruptions during the progression of Alzheimer’s disease, which is consistent with previous literature (Zhi et al., 2024). Similarly, the right inferior temporal gyrus (ITG.R) appears to function as another pivotal hub within the brain network and is selectively disrupted during the progression of Autism Spectrum Disorder (ASD) (Kim et al., 2021).
@@
Furthermore, the high-order organizational patterns identified by HOI-Brain offer novel insights into the diagnosis of brain diseases. Specifically, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease with p-values <0.05<0.05< 0.05 during the progression from cognitively normal (CN) to mild cognitive impairment (MCI). Notably, although these brain regions identified by HOI-Brain demonstrate significant importance in the diagnosis of Alzheimer’s disease (AD) or Autism Spectrum Disorder (ASD) indicating that they may have sustained varying degrees of damage; however, their higher-order interactions exhibit notable differences across various stages of disease progression. This suggests a dissociation between higher-order interactions in the brain and the functions of the individual brain regions themselves, which is consistent with the previous studies (Wang et al., 2019). In addition, we observed that during the transition from healthy individuals to those with Alzheimer’s disease (AD) or autism spectrum disorder (ASD), the positive higher-order interactions between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. On the other hand, the negative higher-order interactions between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s syndrome, which may be due to the fact that Parkinson’s patients, because of their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s syndrome may result from excessive influence of higher-order brain regions on early sensory processing, which is consistent with the enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
@@ 5.2 Limitation and future work
We acknowledge several limitations of our study, which also serve as inspiration for future research: 1)Ignored the discordant signs of negative: In this paper, we focus solely on the concordant signs of positive effects, while only exploring the influence of discordant signs of negative within the interpretable section. These discordant signs are known to play a critical role in the overall system dynamics, as they may reflect redundant or conflicting information across multiple brain regions. Future work should investigate alternative methods to more explicitly address decoherent patterns in the data. 2) Lack of dynamic exploring for higher-order topological features: To mitigate noise in fMRI data, we analyze average statistics across multiple time frames, which helps extract stable low-order and high-order patterns. However, our current approach does not account for the temporal dependencies in higher-order topological features—i.e., how prior time points influence subsequent ones. Since our model reveals that these high-order patterns evolve dynamically over time, further investigation into this temporal evolution could provide novel insights into disease progression. This opens up promising avenues for modeling the dynamic nature of brain network topology. 3) Limited modality of experimental data: Currently, our analysis is based solely on fMRI data. While these data effectively demonstrate the importance of higher-order interactions in brain network analysis, other time-series modalities—such as EEG or MEG—remain unexplored. In future work, we plan to extend our framework to multimodal neuroimaging data and develop methods tailored to capturing higher-order interactions across different temporal scales.
@@ 6 Conclusions
In summary, we propose a new framework, HOI-Brain (Higher-Order Interactions in Brain Network), which can accurately capture signed higher-order interactions in brain networks, extract interpretable signed higher-order topological features, and further exploit the information between lower-order and higher-order features for brain disorder diagnosis. We applied BrainGNN to datasets for Alzheimer’s disease, Parkinson’s syndrome, and autism. With the powerful interpretability, HOI-brain not only performs better on classification than 20 baselines, but also detects salient brain regions associated with classification and discovers important higher-order organizations. Overall, our model shows superiority over alternative graph learning and traditional machine learning classification models. By investigating By analyzing the attention maps of our multi-channel brain transformer, our study identifies salient regions of interest (ROIs) and their key interactions—from a whole-brain perspective—to distinguish brain disorders from healthy controls. Additionally, we uncover higher-order organizational patterns associated with specific disease progression stages. Notably, our framework is generalizable to the analysis of other neuroimaging time-series data. Our proposed framework demonstrates the potential of quadruplet or higher-order region interactions—critical advancements for precision medicine, enhanced understanding of neurological disorders, and broader contributions to neuroimaging research.

### v2

@@ 5.1 The model
Recent conceptual and technical advances in fMRI data analysis provide an opportunity for the field to not only improve the effectiveness of brain disease diagnosis but also deepen our understanding of brain disease mechanisms (Zheng et al., 2024; Luo et al., 2025; Kan et al., 2022b; Yu et al., 2024; Wang et al., 2022; Hao et al., 2023; Bian et al., 2024; Bhattacharya et al., 2025). However, current methods overlook more higher-order patterns with signs, limiting an integrated understanding of brain-wide communication for brain disease diagnosis.
@@
In this study, we focused on the importance of signed higher-order (group) interactions in fMRI signals, which were inferred using a recently developed topological approach (Hyekyoung et al., 2011). Our proposed HOI-Brain incorporates: (i) a novel higher-order interaction representation strategy via Multiplication of Temporal Derivatives (MTD) to quantify dynamic functional co-fluctuations of group ROIs; (ii) a novel higher-order feature extraction method that links higher-order interactions (HOIs) in the brain to higher-order organizational structures, focusing on quadruplet-level interaction signatures and two-dimensional void descriptors extracted through Persistent Homology; (iii) a pioneering effort to distinguish between positively and negatively synergistic higher-order interactions in the brain; (iv) a novel multi-channel brain network Transformer that integrates lower-order edge features with higher-order topological invariants. This new framework emphasizes the potential of quadruplet or higher-order region interactions — representing critical advancements for precision medicine, a deeper understanding of neurological disorders, and broader contributions to neuroimaging research. It demonstrates superior prediction accuracy for classifying Alzheimer’s Disease (AD), Parkinson’s Disease (PD), and Autism Spectrum Disorder (ASD) compared to traditional machine learning models, graph neural network (GNN)-based models, Transformer-based models, hypergraph neural network (HGNN)-based models, and the Persistent Homology (PH)-based models. Additionally, ablation study also indicates that the effectiveness of quadruplet interactions compared to triplet interactions, the effectiveness of distinguishing between positively and negatively synergistic higher-order interactions, and the effectiveness of different components in model.
@@
The findings of our investigation revealed intriguing patterns. In the process of designing an orthonormal clustering readout for the identification of clusters of functionally similar nodes through soft clustering with orthonormal constraints, it was found that the assignment of nodes from different modes to distinct functional modules significantly improves classification accuracy. This finding indicates that both low-level and high-level feature patterns, which are frequently situated within different functional modular structures, are of paramount importance. This observation is in alignment with the conclusions reported in previous studies (Wang et al., 2019). This novel finding also provides a new opportunity for brain disease diagnostic models, specifically how to effectively identify and integrate low-order and higher-order patterns across different functional modules in the brain.
@@
Furthermore, an investigation was conducted into the most discriminative neural patterns using an attention mechanism. This investigation revealed that quadruplet-level interaction signatures exhibit substantially greater importance than two-dimensional void descriptors and edge features in both the ADNI and ABIDE datasets. It is noteworthy that negative synergistic quadruplets tend to exhibit a higher degree of significance in comparison to their positive synergistic counterparts. Furthermore, positive synergistic voids have been observed to exceed negative synergistic voids in terms of their impact. This novel observation may offer critical insights into the pathological mechanisms underlying brain disorders. In order to provide further validation of these findings, inter-group comparisons were performed of signed higher-order topological signatures in the ADNI dataset. The findings demonstrated a substantial discrepancy in the quantity of quadruplet-level interaction signatures among the groups ($p<0.001$) in comparison to void descriptors. Furthermore, positive voids demonstrated more pronounced inter-group disparities than negative voids. The findings suggest that disease progression may primarily disrupt higher-order neural interactions, which subsequently influence high-dimensional hole structures. An intriguing pattern emerged in the comparative analysis of cognitively normal (CN) individuals and those with Alzheimer’s disease (AD). CN individuals appear to rely more on positive information for whole-brain topological organisation, whereas AD patients show a greater dependence on negative information. Furthermore, longitudinal trends indicate that the number of positive quadruplets gradually declines with disease progression, while negative quadruplets follow a biphasic trajectory – decreasing initially and then increasing. This suggests that during the transition from CN to mild cognitive impairment (MCI), there is a decline in both positive and negative synergistic interactions. However, as the condition progresses from MCI to AD, there is an increasing adoption of negative synergy by neural networks, alongside a suppression of excitability. This may be a compensatory mechanism employed by the brain to preserve functional integrity across neural modules.
@@
In the context of investigating the significant brain regions and their interactions in the context of brain diseases, the utilisation of HOI-Brain facilitates the identification of these regions from a more comprehensive perspective, a departure from the low-level perspective that has been adopted in previous studies (Yan et al., 2019; Li et al., 2021). The findings, which function as potential biomarkers, are consistent with a substantial corpus of extant literature (Zidan et al., 2019; Zhi et al., 2024; Ye et al., 2019; Kim et al., 2021). Furthermore, they furnish novel insights with regard to the diagnosis of brain diseases. The right caudate has been identified as a potential hub node within the brain network. Previous literature has indicated that disruptions in this region are consistent with the progression of Alzheimer’s disease (Zhi et al., 2024). In a similar manner, the right inferior temporal gyrus (ITG.R) has been observed to function as an additional pivotal hub within the brain network. It has been demonstrated that this region is selectively disrupted during the progression of Autism Spectrum Disorder (ASD) (Kim et al., 2021).
@@
Furthermore, the higher-order organisational patterns identified by HOI-Brain offer novel insights into the diagnosis of brain diseases. Specifically, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease with p-values <0.05<0.05< 0.05 during the progression from cognitively normal (CN) to mild cognitive impairment (MCI). It is noteworthy that the brain regions identified by HOI-Brain via the analysis of resting-state fMRI data are deemed to be of significant importance in the diagnosis of Alzheimer’s disease (AD) or Autism Spectrum Disorder (ASD). This finding indicates that these brain regions may have sustained varying degrees of damage. However, higher-order interactions exhibited by these regions demonstrate notable differences across various stages of disease progression. This finding is consistent with the conclusions of previous studies (Wang et al., 2019), suggesting a dissociation between higher-order interactions in the brain and the functions of the individual brain regions themselves. Furthermore, observations were made concerning the transition from healthy individuals to those with Alzheimer’s disease (AD) or autism spectrum disorder (ASD). It was noted that the positive higher-order interactions between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. Conversely, the negative higher-order interactions between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with the findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s disease, which may be due to the fact that patients suffering from Parkinson’s disease, due to their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s disease may be attributable to an excessive influence of higher-order brain regions on early sensory processing. This is consistent with enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
@@ 5.2 Limitation and future work
We acknowledge several limitations of our study, which also serve as inspiration for future research: 1) Ignored the discordant signs of negative: In this paper, we focus solely on the concordant signs of positive effects, while only exploring the influence of discordant signs of negative within the interpretable section. These discordant signs are known to play a critical role in the overall system dynamics, as they may reflect redundant or conflicting information across multiple brain regions. Future work should investigate alternative methods to more explicitly address discordant patterns in the data. 2) Lack of dynamic exploring for higher-order topological features: To mitigate noise in fMRI data, we analyze average statistics across multiple time frames, which helps extract stable low-order and higher-order patterns. However, our current approach does not account for the temporal dependencies in higher-order topological features—i.e., how prior time points influence subsequent ones. Since our model reveals that these higher-order patterns evolve dynamically over time, further investigation into this temporal evolution could provide novel insights into disease progression. This opens up promising avenues for modeling the dynamic nature of brain network topology. 3) Limited modality of experimental data: Currently, our analysis is based solely on fMRI data. While these data effectively demonstrate the importance of higher-order interactions in brain network analysis, other time-series modalities—such as EEG or MEG—remain unexplored. In future work, we plan to extend our framework to multimodal neuroimaging data and develop methods tailored to capturing higher-order interactions across different temporal scales.
@@ 6 Conclusions
In summary, we propose a new framework, HOI-Brain (Higher-Order Interactions in Brain Network), which can accurately capture signed higher-order interactions in brain networks, extract interpretable signed higher-order topological features, and further exploit the information between lower-order and higher-order features for brain disorder diagnosis. We applied HOI-Brain to datasets for Alzheimer’s disease, Parkinson’s disease, and Autism Spectrum Disorder. With the powerful interpretability, HOI-brain not only performs better on classification than 20 baselines, but also detects salient brain regions associated with classification and discovers important higher-order organizations. Overall, our model shows superiority over alternative graph learning and traditional machine learning classification models. By analyzing the attention maps of our multi-channel brain Transformer, our study identifies salient regions of interest (ROIs) and their key interactions—from a whole-brain perspective—to distinguish brain disorders from healthy controls. Additionally, we uncover higher-order organizational patterns associated with specific disease progression stages. Notably, our framework is generalizable to the analysis of other neuroimaging time-series data. Our proposed framework demonstrates the potential of quadruplet or higher-order region interactions—critical advancements for precision medicine, enhanced understanding of neurological disorders, and broader contributions to neuroimaging research.